OpenStudy (anonymous):

–4(2x – 1) = –6(x + 2) – 2

OpenStudy (lilg132):

-4(2x - 1) = -6(x + 2) - 2 -8x + 4 = -6x - 12 - 2 - 8x + 6x = - 12 - 2 - 4 -2x = -18 -2x / 2 = -18 / 2 -x = -9 x = 9

OpenStudy (lilg132):

remember to click good answer